OpenStudy (anonymous):

Solve log(3x + 2) = 3. 1000/3 998/3 998 1/3 I think its 998/3

jimthompson5910 (jim_thompson5910):

you are 100% correct

jimthompson5910 (jim_thompson5910):

log(3x + 2) = 3 3x + 2 = 10^3 3x + 2 = 1000 3x = 1000 - 2 3x = 998 x = 998/3

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

Ideally you should check the answer, but it's correct

OpenStudy (anonymous):

thanks (: